Ali walks 220m east, 150m south 40m east. Find the distance and displacement
1
Expert's answer
2020-03-16T13:05:06-0400

The distance


"d=220+150+40=410\:\rm m."

The displacement


"s=\sqrt{(220+40)^2+150^2}=300\:\rm m."